# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v1

@@ Abstract
11footnotetext: The code and data are released at https://github.com/canchen-cc/online-llm-detection
@@ 1 Introduction
While these existing methods can efficiently identify a text source in an offline setting where all the texts to be classified are provided in a single shot, they are not specifically designed to handle scenarios where texts arrive sequentially, and therefore, they might not be directly applicable to the online setting, where certain serious challenges have been observed over the past few years. For example, the American Federal Communications Commission in 2017 decided to repeal net neutrality rules according to the public opinions collected through an online platform (Selyukh, 2017; Weiss, 2019). However, it was ultimately discovered that the overwhelming majority of the total $22$ million comments that support rescinding the rules were machine-generated(Kao, 2017). In 2019, Weiss (2019) used GPT-2 to overwhelm a website for collecting public comments on a medical reform waiver within only four days, where machine-generated comments eventually made up $55.3\%$ of all the comments (more precisely, $1,001$ out of $1,810$ comments). As discussed by Fröhling and Zubiaga (2021), a GPT-J model trained on a politics message board was then deployed on the same forum. It generated posts that included objectionable content and accounted for about $10\%$ of all activity during peak times (Kilcher, 2022). Furthermore, other online attacks mentioned by Fröhling and Zubiaga (2021) may even manipulate public discourse (Ferrara et al., 2016), flood news with fake content (Belz, 2019), or fraud by impersonating others on the Internet or via e-mail (Solaiman et al., 2019). These examples highlight the urgent need for developing algorithms with strong statistical guarantees that can quickly identify machine-generated texts in a timely manner, which, to the best of our knowledge, have been overlooked in the literature.
@@
Our goal, therefore, is to tackle the problem of online detecting LLM-generated texts. More precisely, building upon existing score functions from those “offline approaches”, we aim to quickly determine whether the source of a sequence of texts observed in a streaming fashion is an LLM or a human. Our algorithm leverages the techniques of sequential hypothesis testing (Shafer, 2021; Ramdas et al., 2023; Shekhar and Ramdas, 2023). Specifically, we frame the problem of online detecting LLMs as a sequential hypothesis testing problem, where at each round $t$, a text from an unknown source is observed, and we aim to infer whether it is generated by an LLM. We also assume that a pool of examples of human-written texts is available, and our algorithm can sample a text from this pool of examples at any time $t$. Our method constructs a null hypothesis $H_{0}$ (to be elaborated soon), for which correctly rejecting the null hypothesis implies that the algorithm correctly identifies the source as an LLM under a mild assumption. Furthermore, since it is desirable to quickly identify an LLM when it is present and avoid erroneously declaring the source as an LLM, we also aim to control the type-I error (false positive rate) while maximizing the power to reduce type-II error (false negative rate), and to establish an upper bound on the expected total number of rounds to declare that the source is an LLM. We emphasize that our approach is non-parametric, and hence it does not need to assume that the underlying data of human or machine-generated texts follow a certain distribution (Balsubramani and Ramdas, 2015). It also avoids the need for assuming that the sample size is fixed or to specify it before the testing starts, and hence it is in contrast with some typical hypothesis testing methods that do not enjoy strong statistical guarantees in the anytime-valid fashion (Garson, 2012; Good, 2013; Tartakovsky et al., 2014). The way to achieve these is based on recent developments in sequential hypothesis testing via betting (Shafer, 2021; Shekhar and Ramdas, 2023). We also evaluate the effectiveness of our method through comprehensive experiments. The code and datasets needed to reproduce the experimental results will be made available upon publication.
@@ 2 Preliminaries
$H_{0}\text{ (null hypothesis)}:\mu_{x}=\mu_{y},\quad\text{versus}\quad H_{1}\text{ (alternative hypothesis)}:\mu_{x}\neq\mu_{y}.$ (1)
@@
Sequential Hypothesis Testing by Online Optimization and Betting. Consider the scenario that an online learner engages in multiple rounds of a game with an initial wealth $W_{0}=1$. In each round $t$ of the game, the learner plays a point $\theta_{t}$. Then, the learner receives a fortune after committing $\theta_{t}$, which is $-g_{t}\theta_{t}W_{t-1}$, where $W_{t-1}$ is the learner’s wealth from the previous round $t-1$, and $g_{t}$ can be thought of as “the coin outcome” at $t$ that the learner is trying to “bet” on (Orabona and Pál, 2016). Consequently, the dynamic of the wealth of the learner evolves as:
@@
3:for $t=1,2,\dots,T$ do
@@
12:     Get a hint $d_{t+1}$ and specify the convex decision space $\mathcal{K}_{t+1}:=[-\frac{1}{2d_{t+1}},\frac{1}{2d_{t+1}}]$.
13:     // Update $\theta_{t+1}\in\mathcal{K}_{t+1}$ via $\mathrm{ONS}$ on the loss function $\ell_{t}(\theta):=-\ln(1-g_{t}\theta)$.
14:     Compute $z_{t}=\frac{d\ell_{t}(\theta_{t})}{d\theta}=\frac{g_{t}}{1-g_{t}\theta_{t}}$ and $a_{t}=a_{t-1}+z_{t}^{2}$.
15:     Compute $\theta_{t+1}=\max\left(\min\left(\theta_{t}-\frac{1}{\gamma}\frac{z_{t}}{a_{t}},\frac{1}{2d_{t+1}}\right),-\frac{1}{2d_{t+1}}\right)$.
16:end for
17:if the source has not been declared as an LLM then
18:     Sample $Z\sim\text{Unif}(0,1)$, declare the sequence of texts $y_{t}$ is from an LLM if $W_{T}\geq Z/\alpha$.
19:end if
@@ 3 Our Algorithm
We have covered most of the underlying algorithmic design principles of our online method for detecting LLMs, and we are now ready to introduce our algorithm, which is shown in Algorithm 1. Compared to existing works on sequential hypothesis testing via betting (e.g., Shekhar and Ramdas (2023); Chugg et al. (2023)), which assume knowledge of a bound on the magnitude of the “coin outcome" $g_{t}$ in the learner’s wealth dynamic (2) for all time steps before the testing begins (i.e., assuming prior knowledge of $d_{*}:=\max_{t}|g_{t}|$), we relax this assumption.
Specifically, we consider the scenario where an upper bound on $|g_{t+1}|$, which is denoted by $d_{t+1}$, is available before updating $\theta_{t+1}$ at each round $t$. Our algorithm then plays a point in the decision space $\mathcal{K}_{t+1}$ that guarantees the learner’s wealth remains a non-negative supermartingale (Step 11 in Algorithm 1). We note that if the bound of the output of the underlying score function $\phi(\cdot)$ is known a priori, this scenario holds naturally. Otherwise, we can estimate an upper bound for $|g_{t}|$ for all $t$ based on the first few time steps and execute the algorithm thereafter. One approach is to set the estimate as a conservatively large constant, e.g., twice the maximum value observed in the first few time steps. We observe that this estimate works for our algorithm with most of the score functions $\phi(\cdot)$ that we consider in the experiments. On the other hand, we note that a tighter bound $d_{t}$ will lead to a faster time to reject $H_{0}$ when the unknown source is an LLM, as indicated by the following propositions.
@@
Remark 1. Under the additional assumption of the existence of a good score function $\phi(\cdot)$ that can generate scores with different means for human-written texts and LLM-generated ones, Proposition 1 implies that when the unknown source is declared by Algorithm 1 as an LLM, the probability of this declaration being false will be bounded by $\alpha$. Additionally, if the unknown source is indeed an LLM, then our algorithm can guarantees that it will eventually detect the LLM, since it has asymptotic power one. Moreover, Proposition 1 also provides a non-asymptotic result (see (3)) for bounding the expected time to reject the null hypothesis $H_{0}$, which is also the expected time to declare that the unknown source is an LLM. The bound indicates that a larger difference of the means $\Delta$ can lead to a shorter time to reject the null $H_{0}$.
@@
$H_{0}^{A}:\mu_{x}-\mu_{y}-\epsilon\leq 0\text{ vs.~{}}H_{1}^{A}:\mu_{x}-\mu_{y}-\epsilon>0\text{ and }H_{0}^{B}:\mu_{y}-\mu_{x}-\epsilon\leq 0\text{ vs.~{}}H_{1}^{B}:\mu_{y}-\mu_{x}-\epsilon>0.$ (4)
@@
Remark 2. Proposition 2 indicates that even if there is a difference $\epsilon$ in mean scores between texts written by different humans, the probability that the source is incorrectly declared by Algorithm 3 as an LLM can be controlled below $\alpha$. Besides, our algorithm will eventually declare the source as an LLM if the texts are indeed LLM-generated, as it achieves asymptotic power of one. The expected time to reject $H_{0}$ and then declare the unknown source as an LLM is bounded (see (5)). The bound implies that smaller $\epsilon$ and larger $\Delta$ will result in a shorter time to reject $H_{0}$.
@@ 4.1 Settings
Score Functions. We use $10$ score functions in total from the related works for the experiments. As mentioned earlier, a score function takes a text as an input and outputs a score. For example, one of the configurations of our algorithm that we try uses a score function called Likelihood, which is based on the average of the logarithmic probabilities of each token conditioned on its preceding tokens(Solaiman et al., 2019; Hashimoto et al., 2019). More precisely, for a text $x$ which consists of $n$ tokens, this score function can be formulated as $\phi(x)=\frac{1}{n}\sum_{j=1}^{n}\log p_{\theta}(x_{j}|x_{1:j-1})$, where $x_{j}$ denotes the $j$-th token of the text $x$, $x_{1:j-1}$ means the first $j-1$ tokens, and $p_{\theta}$ represents the probability computed by a language model used for scoring. The score functions considered in the experiments are: 1. DetectGPT: perturbation discrepancy(Mitchell et al., 2023). 2. Fast-DetectGPT: conditional probability curvature(Bao et al., 2023). 3.LRR: likelihood log-rank ratio(Su et al., 2023). 4. NPR: normalized perturbed log-rank(Su et al., 2023). 5. DNA-GPT: WScore(Yang et al., 2023). 6. Likelihood: mean log probabilities (Solaiman et al., 2019; Hashimoto et al., 2019; Gehrmann et al., 2019). 7. LogRank: averaged log-rank in descending order by probabilities(Gehrmann et al., 2019). 8. Entropy: mean token entropy of the predictive distribution(Gehrmann et al., 2019; Solaiman et al., 2019; Ippolito et al., 2019). 9. RoBERTa-base: a pre-trained classifier(Liu et al., 2019). 10. RoBERTa-large: a larger pre-trained classifier with more layers and parameters(Liu et al., 2019). The first eight score functions calculate scores based on certain statistical properties of texts, with each text’s score computed via a language model. The last two score functions compute scores by using some pre-trained classifiers. For the reader’s convenience, more details about the implementation of the score functions $\phi(\cdot)$ are provided in Appendix B.
LLMs and Datasets. Our experiments focus on the black-box setting(Bao et al., 2023), which means that if $x$ is generated by a model $q_{s}$, i.e., $x\sim q_{s}$, a different model $p_{\theta}$ will then be used to evaluate the metrics such as the log-probability $\log p_{\theta}(x)$ when calculating $\phi(x)$. The models $q_{s}$ and $p_{\theta}$ are respectively called the “source model" and “scoring model" for clarity. The black-box setting is a relevant scenario in practice because the source model used for generating the texts to be inferred is likely unknown in practice, which makes it elusive to use the same model to compute the scores.
@@
We construct a dataset that contains some real news and fake ones generated by LLMs for 2024 Olympics. Specifically, we collect 500 news about Paris 2024 Olympic Games from its official website(Olympics, 2024) and then use three source models, Gemini-1.5-Flash, Gemini-1.5-Pro(Google Cloud, 2024a), and PaLM 2(Google Cloud, 2024b; Chowdhery et al., 2023) to generate an equal number of fake news based on the first 30 tokens of each real one respectively. Two scoring models for computing the text scores $\phi(\cdot)$ are considered, which are GPT-Neo-2.7B (Neo-2.7)(Black et al., 2021) and Gemma-2B(Google, 2024). The perturbation model that is required for the score function DetectGPT and NPR is T5-3B(Raffel et al., 2020). For Fast-DetectGPT, the sampling model is GPT-J-6B(Wang and Komatsuzaki, 2021) when scored with Neo-2.7, and Gemma-2B when the scoring model is Gemma-2B. We sample human-written text $x_{t}$ from a pool of 500 news articles from the XSum dataset(Narayan et al., 2018). We also consider existing datasets from Bao et al. (2023) for the experiments. Details can be found in Appendix H.
@@
Baselines. Our method are compared with two baselines, which adapt the fixed-time permutation test (Good, 2013) to the scenario of the sequential hypothesis testing. Specifically, the first baseline conducts a permutation test after collecting every $k$ samples. If the result does not reject $H_{0}$, then it will wait and collect another $k$ samples to conduct another permutation test on this new batch of $k$ samples. This process is repeated until $H_{0}$ is rejected or the time $t$ runs out (i.e., when $t=T$). The significance leave parameter of the permutation test is set to be the same constant $\alpha$ for each batch, which does not maintain a valid level-$\alpha$ test overall. The second baseline is similar to the first one except that the significance level parameter for the $i$-th batch is set to be $\alpha/2^{i}$, with $i$ starting from $1$, which aims to ensure that the cumulative type-I error is bounded by $\alpha$ via the union bound. The detailed process of the baselines is described in Appendix G.
@@
(a) Averaged test results with text $x_{t}$ sampled from XSum and $y_{t}$ from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Neo-2.7.
(b) Averaged test results with text $x_{t}$ sampled from XSum and $y_{t}$ from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Gemma-2B.
4.2 Experimental Results
The experiments evaluate the performance of our method and baselines under both $H_{0}$ and $H_{1}$. As there is inherent randomness from the observed samples of the texts in the online setting, we repeat $1000$ runs and report the average results over these $1000$ runs. Specifically, we report the false positive rate (FPR) under $H_{0}$, which is the number of times the source of $y_{t}$ is incorrectly declared as an LLM when it is actually human, divided by the total of $1000$ runs. We also report the average time to reject the null under $H_{1}$ (denoted as Rejection Time $\tau$), which is the time our algorithm takes to reject $H_{0}$ and correctly identify the source when $y_{t}$ is indeed generated by an LLM. More precisely, the rejection time $\tau$ is the average time at which either $W_{t}^{A}$ or $W_{t}^{B}$ exceeds $2/\alpha$ before $T$; otherwise, $\tau$ is set to $T=500$, regardless of whether it rejects $H_{0}$ at $T$, since the the time budget runs out. The parameter value $d_{t}\leftarrow d_{*}$ in Scenario 1 (oracle) is shown in Table 2, and the value for $\epsilon$ can be found in Table 1 in the appendix. For the estimated $\epsilon$ and $d_{t}$ of each sequential testing in Scenario 2, they are displayed in Table 5 and Table 6 respectively in the appendix. Our method and the baselines require specifying the significance level parameter $\alpha$. In our experiments, we try $20$ evenly spaced values of the significance level parameter $\alpha$ that ranges from $0.005$ to $0.1$ and report the performance of each one.
@@ 4.2 Experimental Results
(a) Averaged test results with text $x_{t}$ sampled from XSum and $y_{t}$ from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Neo-2.7.
(b) Averaged test results with text $x_{t}$ sampled from XSum and $y_{t}$ from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Gemma-2B.
@@
Figure 3 shows the empirical results of our algorithm under Scenario 2, where it has to use the first few samples to specify $d_{t}$ and $\epsilon$ before starting the algorithm. Under this scenario, our algorithm equipped with most of the score functions still perform effectively. We observe that our algorithm with 1) Fast-DetectGPT as the score function $\phi(\cdot)$ and Neo-2.7 as the language model for computing the score, and with 2) Likelihood as the score function $\phi(\cdot)$ and Gemma-2B for computing the value of $\phi(\cdot)$ have the best performance under this scenario. Compared to the first case where the oracle of $d_{*}$ and $\epsilon$ is available and exploited, these two configurations only result in a slight degradation of the performance under Scenario 2, and we note that our algorithm can only start updating after the first $10$ time steps under this scenario. We observe that the bound of $d_{*}$ that we estimated using the samples collected from first $10$ time steps is significantly larger than the tightest bound of $d_{*}$ in most of the runs where we refer the reader to Table 2, 6 in the appendix for details, which explains why most of the configurations under Scenario 2 need a longer time to detect LLMs, as predicted by our propositions. We also observe that the configurations with two supervised classifiers (RoBERTa-based and RoBERTa-large) and the combinations of a couple of score functions and the scoring model Gemma-2B do not strictly control FPRs across all significance levels. This is because the estimated $d_{t}$ for these score functions is not large enough to ensure that the wealth $W_{t}$ remains nonnegative at all time points $t$. That is, we observed $2d_{t}<\lvert\phi(x_{t})-\phi(y_{t})\rvert+\epsilon$ for some $t$ in the experiments, and hence the wealth $W_{t}$ is no longer a non-negative supermartingale, which prevents the application of Ville’s inequality to guarantee a level-$\alpha$ test. Nevertheless, our algorithm with eight score functions that utilize the scoring model Neo-2.7 can still effectively control type-I error and detect LLMs by around $t=300$.

### v2

@@ Abstract
11footnotetext: The code and data are released at https://github.com/canchen-cc/online-llm-detection.
@@ 1 Introduction
While these existing methods can efficiently identify a text source in an offline setting where all the texts to be classified are provided in a single shot, they are not specifically designed to handle scenarios where texts arrive sequentially, and the goal is to quickly detect if the source is an LLM with strong statistical guarantees. Moreover, the process of most of the related works for offline detecting LLMs involves computing a score for each sample to determine how likely it is machine-generated and classify a text as a machine-generated one by determining whether the score passes a certain threshold. The threshold is a critical parameter that needs to be tuned using a validation dataset and must be determined before classifying testing data. Therefore, these offline detection methods fall short in the online scenarios, where one would like to design efficient algorithms with robust performance metrics, such as being a valid level-$\alpha$ test, having asymptotic power $1$, and enjoying a bound on the expected detection time for LLMs, which can be appealing for time-sensitive applications like monitoring online platforms.
@@
For example, the American Federal Communications Commission in 2017 decided to repeal net neutrality rules according to the public opinions collected through an online platform (Selyukh, 2017; Weiss, 2019). However, it was ultimately discovered that the overwhelming majority of the total $22$ million comments that support rescinding the rules were machine-generated (Kao, 2017). In 2019, Weiss (2019) used GPT-2 to overwhelm a website for collecting public comments on a medical reform waiver within only four days, where machine-generated comments eventually made up $55.3\%$ of all the comments (more precisely, $1,001$ out of $1,810$ comments). As discussed by Fröhling and Zubiaga (2021), a GPT-J model trained on a politics message board was then deployed on the same forum. It generated posts that included objectionable content and accounted for about $10\%$ of all activity during peak times (Kilcher, 2022). Furthermore, other online attacks mentioned by Fröhling and Zubiaga (2021) may even manipulate public discourse (Ferrara et al., 2016), flood news with fake content (Belz, 2019), or fraud by impersonating others on the Internet or via e-mail (Solaiman et al., 2019). However, existing bot detection methods for social media (e.g., Davis et al. (2016); Varol et al. (2017); Pozzana and Ferrara (2020); Ferrara (2023) and the references therein) might not be directly applicable to the online setting with strong statistical guarantees, to the best of our knowledge, and they often require training on extensive labeled datasets beforehand. This highlights the urgent need to develop algorithms with strong statistical guarantees that can quickly and effectively identify machine-generated texts in a timely manner, which has been largely overlooked in the literature.
@@
Our goal, therefore, is to tackle the problem of online detecting LLM-generated texts. More precisely, building upon existing score functions from those “offline approaches”, we aim to quickly determine whether the source of a sequence of texts observed in a streaming fashion is an LLM or a human, with a concrete bound on the required samples to detect LLMs and without the need for tuning a threshold of the affinity score that is typically used to classify data in those offline approaches. Our algorithm leverages the techniques of sequential hypothesis testing (Shafer, 2021; Ramdas et al., 2023; Shekhar and Ramdas, 2023). Specifically, we frame the problem of online detecting LLMs as a sequential hypothesis testing problem, where at each round $t$, a text from an unknown source is observed, and we aim to infer whether it is generated by an LLM. We also assume that a pool of examples of human-written texts is available, and our algorithm can sample a text from this pool of examples at any time $t$. Our method constructs a null hypothesis $H_{0}$ (to be elaborated soon), for which correctly rejecting the null hypothesis implies that the algorithm correctly identifies the source as an LLM under a mild assumption. Furthermore, since it is desirable to quickly identify an LLM when it is present and avoid erroneously declaring the source as an LLM, we also aim to control the type-I error (false positive rate) while maximizing the power to reduce type-II error (false negative rate), and to establish an upper bound on the expected total number of rounds to declare that the source is an LLM. We emphasize that our approach is non-parametric, and hence it does not need to assume that the underlying data of human or machine-generated texts follow a certain distribution (Balsubramani and Ramdas, 2015). It also avoids the need for assuming that the sample size is fixed or to specify it before the testing starts, and hence it is in contrast with some typical hypothesis testing methods that do not enjoy strong statistical guarantees in the anytime-valid fashion (Garson, 2012; Good, 2013; Tartakovsky et al., 2014). The way to achieve these is based on recent developments in sequential hypothesis testing via betting (Shafer, 2021; Shekhar and Ramdas, 2023). The setting of online testing with anytime-valid guarantees could be particularly useful when one seeks substantial savings in both data collection and time without compromising the reliability of their statistical testing. These desiderata might be elusive for approaches that based on collecting data in batch and classifying them offline to achieve. The effectiveness of our method is evaluated through comprehensive experiments.
@@ 2 Preliminaries
[…]
@@
Sequential Hypothesis Testing by Betting. Consider the scenario that an online learner engages in multiple rounds of a game with an initial wealth $W_{0}=1$. In each round $t$ of the game, the learner plays a point $\theta_{t}$. Then, the learner receives a fortune after committing $\theta_{t}$, which is $-g_{t}\theta_{t}W_{t-1}$, where $W_{t-1}$ is the learner’s wealth from the previous round $t-1$, and $g_{t}$ can be thought of as “the coin outcome” at $t$ that the learner is trying to “bet” on (Orabona and Pál, 2016). Consequently, the dynamic of the wealth of the learner evolves as:
@@
3:# $T$ is the time budget, which can be $\infty$ if their is no time constraint.
4:for $t=1,2,\dots,T$ do
@@
12:     Get a hint $d_{t+1}$ which satisfies $d_{t+1}\geq|g_{t+1}|$.
13:     Specify the decision space $\mathcal{K}_{t+1}:=[-\frac{1}{2d_{t+1}},\frac{1}{2d_{t+1}}]$ to ensure $W_{t+1}$ is nonnegative.
14:     // Update $\theta_{t+1}\in\mathcal{K}_{t+1}$ via $\mathrm{ONS}$ on the loss function $\ell_{t}(\theta):=-\ln(1-g_{t}\theta)$.
15:     Compute $z_{t}=\frac{d\ell_{t}(\theta_{t})}{d\theta}=\frac{g_{t}}{1-g_{t}\theta_{t}}$ and $a_{t}=a_{t-1}+z_{t}^{2}$.
16:     Compute $\theta_{t+1}=\max\left(\min\left(\theta_{t}-\frac{1}{\gamma}\frac{z_{t}}{a_{t}},\frac{1}{2d_{t+1}}\right),-\frac{1}{2d_{t+1}}\right)$.
17:end for
18:if the source has not been declared as an LLM then
19:     Sample $Z\sim\text{Unif}(0,1)$, declare the sequence of texts $y_{t}$ is from an LLM if $W_{T}\geq Z/\alpha$.
20:end if
@@ 3 Our Algorithm
We have covered most of the underlying algorithmic design principles of our online method for detecting LLMs, and we are now ready to introduce our algorithm, which is shown in Algorithm 1. Compared to existing works on sequential hypothesis testing via betting (e.g., Shekhar and Ramdas (2023); Chugg et al. (2023)), which assume knowledge of a bound on the magnitude of the “coin outcome" $g_{t}$ in the learner’s wealth dynamic (1) for all time steps before the testing begins (e.g., assuming for all $t$, $\phi(x_{t})$ and $\phi(y_{t})$ are always in $[0,1]$), we relax this assumption, which is motivated by the concern that a good estimate of the bound for the coin outcome might not be available before the sequential testing starts. Specifically, we consider the scenario where an upper bound on $|g_{t+1}|$ at round $t+1$, which is denoted by $d_{t+1}$, is available before updating $\theta_{t+1}$ at each round $t$. Our algorithm then plays a point in the decision space $\mathcal{K}_{t+1}$ that guarantees the learner’s wealth remains a non-negative supermartingale (Step 15 in Algorithm 1). We note that if the bound of the output of the underlying score function $\phi(\cdot)$ is known a priori, this scenario holds naturally. Otherwise, we can estimate an upper bound for $|g_{t}|$ for all $t$ based on the first few time steps and execute the algorithm thereafter. One approach is to set the estimate as a conservatively large constant, e.g., twice the maximum value observed in the first few time steps. We observe that this estimate works for our algorithm with most of the score functions $\phi(\cdot)$ that we consider in the experiments. On the other hand, we note that a tighter bound $d_{t}$ will lead to a faster time to reject $H_{0}$ when the unknown source is an LLM, as indicated by the following propositions.
@@
Remark 1. Under the additional assumption of the existence of a good score function $\phi(\cdot)$ that can generate scores with different means for human-written texts and LLM-generated ones, Proposition 1 implies that when the unknown source is declared by Algorithm 1 as an LLM, the probability of this declaration being false will be bounded by $\alpha$. Additionally, if the unknown source is indeed an LLM, then our algorithm can guarantee that it will eventually detect the LLM, since it has asymptotic power one. Moreover, Proposition 1 also provides a non-asymptotic result for bounding the expected time to reject the null hypothesis $H_{0}$, which is also the expected time to declare that the unknown source is an LLM. The bound indicates that a larger difference of the means $\Delta$ can lead to a shorter time to reject the null $H_{0}$.
@@
$H_{0}^{A}:\mu_{x}-\mu_{y}-\epsilon\leq 0\text{ vs.~{}}H_{1}^{A}:\mu_{x}-\mu_{y}-\epsilon>0,$
$H_{0}^{B}:\mu_{y}-\mu_{x}-\epsilon\leq 0\text{ vs.~{}}H_{1}^{B}:\mu_{y}-\mu_{x}-\epsilon>0.$
@@
Remark 2. Proposition 2 indicates that even if there is a difference $\epsilon$ in mean scores between texts written by different humans, the probability that the source is incorrectly declared by Algorithm 3 as an LLM can be controlled below $\alpha$. Furthermore, the bound on $\mathbb{E}[\tau]$ implies that smaller $\epsilon$ and larger $\Delta$ will result in a shorter time to reject $H_{0}$.
@@ 4 Experiments
Score Functions. We use $10$ score functions in total from the related works for the experiments. As mentioned earlier, a score function takes a text as an input and outputs a score. For example, one of the configurations of our algorithm that we try uses a score function called Likelihood, which is based on the average of the logarithmic probabilities of each token conditioned on its preceding tokens (Solaiman et al., 2019; Hashimoto et al., 2019). More precisely, for a text $x$ which consists of $n$ tokens, this score function can be formulated as $\phi(x)=\frac{1}{n}\sum_{j=1}^{n}\log p_{\theta}(x_{j}|x_{1:j-1})$, where $x_{j}$ denotes the $j$-th token of the text $x$, $x_{1:j-1}$ means the first $j-1$ tokens, and $p_{\theta}$ represents the probability computed by a language model used for scoring. The score functions that we considered in the experiments are: 1. DetectGPT: perturbation discrepancy (Mitchell et al., 2023). 2. Fast-DetectGPT: conditional probability curvature (Bao et al., 2023). 3.LRR: likelihood log-rank ratio (Su et al., 2023). 4. NPR: normalized perturbed log-rank (Su et al., 2023). 5. DNA-GPT: WScore (Yang et al., 2023). 6. Likelihood: mean log probabilities (Solaiman et al., 2019; Hashimoto et al., 2019; Gehrmann et al., 2019). 7. LogRank: averaged log-rank in descending order by probabilities (Gehrmann et al., 2019). 8. Entropy: mean token entropy of the predictive distribution (Gehrmann et al., 2019; Solaiman et al., 2019; Ippolito et al., 2019). 9. RoBERTa-base: a pre-trained classifier (Liu et al., 2019). 10. RoBERTa-large: a larger pre-trained classifier with more layers and parameters (Liu et al., 2019). The first eight score functions calculate scores based on certain statistical properties of texts, with each text’s score computed via a language model. The last two score functions compute scores by using some pre-trained classifiers. For the reader’s convenience, more details about the implementation of the score functions $\phi(\cdot)$ are provided in Appendix B.
@@
LLMs and Datasets. Our experiments focus on the black-box setting (Bao et al., 2023), which means that if $x$ is generated by a model $q_{s}$, i.e., $x\sim q_{s}$, a different model $p_{\theta}$ will then be used to evaluate the metrics such as the log-probability $\log p_{\theta}(x)$ when calculating $\phi(x)$. The models $q_{s}$ and $p_{\theta}$ are respectively called the “source model" and “scoring model" for clarity. The black-box setting is a relevant scenario in practice because the source model used for generating the texts to be inferred is likely unknown in practice, which makes it elusive to use the same model to compute the scores. We construct a dataset that contains some real news and fake ones generated by LLMs for 2024 Olympics. Specifically, we collect 500 news about Paris 2024 Olympic Games from its official website (Olympics, 2024) and then use three source models, Gemini-1.5-Flash, Gemini-1.5-Pro (Google Cloud, 2024a), and PaLM 2 (Google Cloud, 2024b; Chowdhery et al., 2023) to generate an equal number of fake news based on the first 30 tokens of each real one respectively. Two scoring models for computing the text scores $\phi(\cdot)$ are considered, which are GPT-Neo-2.7B (Neo-2.7) (Black et al., 2021) and Gemma-2B (Google, 2024). The perturbation model that is required for the score function DetectGPT and NPR is T5-3B (Raffel et al., 2020). For Fast-DetectGPT, the sampling model is GPT-J-6B (Wang and Komatsuzaki, 2021) when scored with Neo-2.7, and Gemma-2B when the scoring model is Gemma-2B. We sample human-written text $x_{t}$ from a pool of 500 news articles from the XSum dataset (Narayan et al., 2018). We emphasize that we also consider existing datasets from Bao et al. (2023) for the experiments. Details can be found in Appendix H, which also demonstrate the effectiveness of our method.
@@
Baselines. Our method are compared with two baselines, which adapt the fixed-time permutation test  (Good, 2013) to the scenario of the sequential hypothesis testing. Specifically, the first baseline conducts a permutation test after collecting every $k$ samples. If the result does not reject $H_{0}$, then it will wait and collect another $k$ samples to conduct another permutation test on this new batch of $k$ samples. This process is repeated until $H_{0}$ is rejected or the time $t$ runs out (i.e., when $t=T$). The significance level parameter of the permutation test is set to be the same constant $\alpha$ for each batch, which does not maintain a valid level-$\alpha$ test overall. The second baseline is similar to the first one except that the significance level parameter for the $i$-th batch is set to be $\alpha/2^{i}$, with $i$ starting from $1$, which aims to ensure that the cumulative type-I error is bounded by $\alpha$ via the union bound. The detailed process of the baselines is described in Appendix G.
(a) Averaged results with text $x_{t}$ sampled from XSum and $y_{t}$ from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Neo-2.7.
(b) Averaged results with text $x_{t}$ sampled from XSum and $y_{t}$ from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Gemma-2B.
@@
Experimental Results. The experiments evaluate the performance of our method and baselines under both $H_{0}$ and $H_{1}$. As there is inherent randomness from the observed samples of the texts in the online setting, we repeat $1000$ runs and report the average results over these $1000$ runs. Specifically, we report the false positive rate (FPR) under $H_{0}$, which is the number of times the source of $y_{t}$ is incorrectly declared as an LLM when it is actually human, divided by the total of $1000$ runs. We also report the average time to reject the null under $H_{1}$ (denoted as Rejection Time $\tau$), which is the time our algorithm takes to reject $H_{0}$ and correctly identify the source when $y_{t}$ is indeed generated by an LLM. More precisely, the rejection time $\tau$ is the average time at which either $W_{t}^{A}$ or $W_{t}^{B}$ exceeds $2/\alpha$ before $T$; otherwise, $\tau$ is set to $T=500$, regardless of whether it rejects $H_{0}$ at $T$, since the the time budget runs out. The parameter value $d_{t}\leftarrow d_{*}$ in Scenario 1 (oracle) is shown in Table 2, and the value for $\epsilon$ can be found in Table 1 in the appendix. For the estimated $\epsilon$ and $d_{t}$ of each sequential testing in Scenario 2, they are displayed in Table 5 and Table 6 respectively in the appendix. Our method and the baselines require specifying the significance level parameter $\alpha$. In our experiments, we try $20$ evenly spaced values of the significance level parameter $\alpha$ that ranges from $0.005$ to $0.1$ and report the performance of each one.
@@
(a) Averaged results with text $x_{t}$ sampled from XSum and $y_{t}$ from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Neo-2.7.
(b) Averaged results with text $x_{t}$ sampled from XSum and $y_{t}$ from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Gemma-2B.
@@
Figure 3 shows the empirical results of our algorithm under Scenario 2, where it has to use the first few samples to specify $d_{t}$ and $\epsilon$ before starting the algorithm. Under this scenario, our algorithm equipped with most of the score functions still performs effectively. We observe that our algorithm with 1) Fast-DetectGPT as the score function $\phi(\cdot)$ and Neo-2.7 as the language model for computing the score, and with 2) Likelihood as the score function $\phi(\cdot)$ and Gemma-2B for computing the value of $\phi(\cdot)$ have the best performance under this scenario. Compared to the first case where the oracle of $d_{*}$ and $\epsilon$ is available and exploited, these two configurations only result in a slight degradation of the performance under Scenario 2, and we note that our algorithm can only start updating after the first $10$ time steps under this scenario. We observe that the bound of $d_{*}$ that we estimated using the samples collected from first $10$ time steps is significantly larger than the tightest bound of $d_{*}$ in most of the runs where we refer the reader to Table 2, 6 in the appendix for details, which explains why most of the configurations under Scenario 2 need a longer time to detect LLMs, as predicted by our propositions. We also observe that the configurations with two supervised classifiers (RoBERTa-based and RoBERTa-large) and the combinations of a couple of score functions and the scoring model Gemma-2B do not strictly control FPRs across all significance levels. This is because the estimated $d_{t}$ for these score functions is not large enough to ensure that the wealth $W_{t}$ remains nonnegative at all time points $t$. That is, we observed $2d_{t}<\lvert\phi(x_{t})-\phi(y_{t})\rvert+\epsilon$ for some $t$ in the experiments, and hence the wealth $W_{t}$ is no longer a non-negative supermartingale, which prevents the application of Ville’s inequality to guarantee a level-$\alpha$ test. Nevertheless, our algorithm with eight score functions that utilize the scoring model Neo-2.7 can still effectively control type-I error and detect LLMs by around $t=300$.